1.Sallie was able to buy a blouse at 35% discount. If the original price is Php 850, how much discount was she able to avail? How much is the fiscounted price?
2. For being the top 2 of her batch, patricia was offered a 25% discount on her tuition fee for the following semester. If she only paid Php 51,000, how much is her tuition fee?
3. In a recent sale, michelle was able to buy a microwave oven for Php 12,500. Of the original price is Php 15,625, what is the rate of discount?
4. Rickson was able to sell a house and lot for Php 4,500,00. If he earned a commission of Php 157,500. What is the rate of commission?
5. Arlene, a sales executive, is given an 8% monthly commission of all sales for s month in excess of Php 500,000. Of her total sales for the month of march is Php 612,800, how much commission will she receive for the month of may?
1
Expert's answer
2021-10-04T18:33:27-0400

1) discount availed "=35/100\times 850=Php297.5"

discounted price "850-297.5=Php552.5"

2) Php51000 represents 75%

what about 100%

tuition fee "=100/75\times 51000=Php68000"

3) "discount=15625-12500=Php3125"

rate of discount "=3125/15625\times 100\%=20\%"

4) rate of commission ="157500/450000\times 100\%=35\%"

5) commission =8%

excess sales=612800-50000=112800

commission received = "8/100\times 112800=Php9024"